what is dot product

Grade:12

3 Answers

raju
10 years ago
In mathematics, the dot product, or scalar product (or sometimes inner product in the context of Euclidean space), is an algebraic operation that takes two equal-length sequences of numbers (usually coordinate vectors) and returns a single number. This operation can be defined either algebraically or geometrically. Algebraically, it is the sum of the products of the corresponding entries of the two sequences of numbers. Geometrically, it is the product of the magnitudes of the two vectors and the cosine of the angle between them. The name "dot product" is derived from the centered dot " · " that is often used to designate this operation; the alternative name "scalar product" emphasizes the scalar (rather than vectorial) nature of the result. In three-dimensional space, the dot product contrasts with the cross product of two vectors, which produces a pseudovector as the result. The dot product is directly related to the cosine of the angle between two vectors in Euclidean space of any number of dimensions.
Sher Mohammad IIT Delhi
10 years ago
let two vectors, (x1i+x2j+x3k) and(y1i+y2j+y3k) , there dot product is x1y1+x2y2+x3y3
example:(1i+2j+3k) ,(2i+2j+4k)has dot product 2+2*2+3*4=18

Sher Mohammad
B.Tech, IIT Delhi

Indu
10 years ago
A · B = |A| * |B| * cos(T) A · B = 1 * 1 * cos(T) A · B = cos(T) In this case, the dot product is equal to the cosine of the angle between the vectors.